OpenStudy (anonymous):

verify the identity 2(tan x-cot x)/tan^2 x-cot^2 x=sin 2x

OpenStudy (experimentx):

2(tan x-cot x)/tan^2 x-cot^2 x = 2(tan x-cot x)/(tan x-cot x)(tan x+cot x) = 2/(tan x+cot x) = 2/(sin x/ cos x + cos x/ sinx) = 2/(sin ^2 x + cos ^2 x)/sin x cos x = 2sin x cos x/(sin ^2 x + cos ^2 x) = 2sin x cos x/(1) = sin 2x

OpenStudy (anonymous):

(tan x-cot x) cancels right?

OpenStudy (experimentx):

yes

OpenStudy (anonymous):

ohhh it cancels after you expanded the bottom

OpenStudy (experimentx):

yes

OpenStudy (anonymous):

im terrible at this concept, thanks so much

OpenStudy (experimentx):

you are welcome